# Kernel Ridge Riesz Representers: Generalization Error and Mis-specification

### v3

@@ 6 Proof of main result via the counterfactual effective dimension
The steps are as follows: (i) translate the desired statement from generalization error to $H$ norm (Lemma 6.1); (ii) characterize high probability events via concentration in $H$ (Proposition 6.1); (iii) use these high probability events to prove a nonasymptotic bound in terms of standard learning theory quantities (Proposition 6.2).
Among the high probability events is one that involves the counterfactual features $\phi^{(m)}(W)$. It pertains to the variance of the KRRR estimator $\tilde{\alpha}$. To justify this high probability event, I bound the counterfactual effective dimension of $\phi^{(m)}(W)$ in terms of the actual effective dimension of $\phi(W)$ (Lemma 6.2).
The main result (Theorem 5.1) then follows from simplifying the nonasymptotic bound (Proposition 6.2) using known bounds on the learning theory quantities that hold under smoothness (Assumption 3.2) and spectral decay (Assumption 3.3) conditions. The latter controls the actual effective dimension. I defer proofs of lemmas to Appendix D.
Lemma 6.1 (From generalization error to $H$ norm; c.f. Lemma 12 of [Fischer and Steinwart, 2020]).
Suppose the conditions of Lemma 4.1 hold. Then $\mathcal{R}(\tilde{\alpha})=\|T^{\frac{1}{2}}(\tilde{\alpha}-\alpha_{H})\|^{2}_{H}.$
Definition 6.1 (Standard learning theory quantities).
@@
Definition 6.2 (New learning theory quantity).
Define the counterfactual effective dimension as $\mathcal{N}^{(m)}(\lambda)=Tr\{(T+\lambda)^{-1}T^{(m)}\}$, where $T^{(m)}=\mathbb{E}[\phi^{(m)}(W)\otimes\{\phi^{(m)}(W)\}^{*}]$ is the counterfactual covariance operator.
Lemma 6.2 (New learning theory bound).
Suppose the conditions of Lemma 4.1 hold. Then $\mathcal{N}^{(m)}(\lambda)\leq\bar{M}\cdot\mathcal{N}(\lambda).$
Remark 6.1 (Avoiding an auxiliary approximation assumption).
Lemma 6.2 demonstrates, by a simple argument, that a standard assumption in semiparametric theory (Assumption 3.1) implies that the counterfactual effective dimension (Definition 6.2) is not much greater than the actual effective dimension (Definition 6.1). This connection between semiparametric theory and learning theory appears to be new, and it powers the main result.
Previous works that study generalization error, for different estimators of Riesz representers, appear to place auxiliary approximation assumptions to limit the counterfactual effective dimension, e.g. [Chernozhukov et al., 2022d, Assumption SC(a)], [Chernozhukov et al., 2022c, Assumption 2], and [Chernozhukov et al., 2020, Assumption 2].
Definition 6.3 (High probability events in $H$).
@@
Proposition 6.1 (High probability events in $H$).
Suppose the conditions of Lemma 4.1 hold. Then $\mathbb{P}(\mathcal{E}_{j}^{c})\leq\delta/3$ for each $j\in\{1,2,3\}.$
The proof of Proposition 6.1 appeals to Lemma 6.2 in order to handle $\mathcal{E}_{3}$, which contains the counterfactual features via the counterfactual mean embedding $\hat{\mu}^{(m)}=\mathbb{E}_{n}\{\phi^{(m)}(W)\}$.
Proposition 6.2 (Nonasymptotic bound).
Suppose the conditions of Lemma 4.1 hold. If $n$ is sufficiently large that $n\geq 192\ln^{2}(6/\delta)\kappa^{\prime}\mathcal{N}(\lambda)\lambda^{-1}$ and $\lambda\leq\|T\|_{op},$ then with probability $1-\delta$,
$\|T^{1/2}(\tilde{\alpha}-\alpha_{H})\|^{2}_{H}\leq 96\ln^{2}(6/\delta)\mathopen{}\mathclose{{}\left\{\mathcal{A}(\lambda)+\frac{\kappa^{2}\mathcal{B}(\lambda)}{n^{2}\lambda}+\frac{\kappa\mathcal{A}(\lambda)}{n\lambda}+\frac{\kappa^{\prime}\Upsilon^{2}}{n^{2}\lambda}+\frac{\Sigma^{2}\mathcal{N}(\lambda)}{n}}\right\}.$
The proof of Proposition 6.2 appeals to Proposition 6.1 and the union bound.
@@ 7 Simulated and real data analysis
7.1 Nominal pointwise coverage of heterogeneous treatment effects
I present coverage simulations for heterogeneous treatment effects (Example 3.2) viewed as a causal function. To lighten notation, I let $V=X_{1}$ and I write the heterogeneous treatment effects as $\textsc{cate}(v)$.
@@ 7.1 Nominal pointwise coverage of heterogeneous treatment effects
Figure 1: Simulation design of [Abrevaya et al., 2015]. cate(v) is the curve and the three values of $\theta_{0}$ are the three points. The horizontal dashed line is at zero.
I evaluate confidence intervals constructed from the KRRR estimator $\tilde{\alpha}$ (Definition 4.2) in the context of debased machine learning $(\hat{\theta},\hat{\sigma}^{2})$ (Definition 5.1). I consider three variations of $(\hat{\theta},\hat{\sigma}^{2})$, corresponding to different choices of the nonparametric regression estimator $\hat{\gamma}$: lasso, random forest, and neural network. Previous works on classical kernel ridge balancing weights typically disallow $\gamma_{0}\not\in H$, and appear to lack formal inference for heterogeneous treatment effects. This use of kernel ridge balancing weights is justified by the main result of this paper.
Following [Chernozhukov et al., 2023, Tables S1 to S6], I consider low dimensional and high dimensional specifications. In the low dimensional specification, the estimator $\hat{\gamma}$ uses $(D,X)$ as well as their interactions. In the high dimensional specification, $\hat{\gamma}$ uses fourth order polynomials.
Throughout, KRRR uses the Gaussian kernel for $X$, which corresponds to using all Hermite polynomials, appropriately weighted [Rasmussen and Williams, 2006, Section 4.3]. See Appendix E for complete implementation details.
An important choice is how to tune the bandwidth $h$ in $\textsc{local}\{(v-v^{*})/h\}$. I follow the heuristic $h=c_{h}\hat{\sigma}_{v}n^{-1/5}$ [Colangelo and Lee, 2020], where $\hat{\sigma}^{2}_{v}$ is the sample variance of $V$. This heuristic satisfies the theoretical requirements $h=o(1)$ and $n^{-1/2}h^{-3/2}=o(1)$, and it implies a nonparametric rate of convergence of $(nh)^{-1/2}=n^{-2/5}$. The hyperparameter $c_{h}$ is chosen by the analyst. To evaluate robustness to tuning, I consider $c_{h}\in\{0.25,0.50,1.00\}$ below.
Table 1: Low dimensional coverage simulations
Lasso Random Forest Neural Network
Ave. Ave. 95% Ave. Ave. 95% Ave. Ave. 95%
$v^{*}$ cate($v^{*}$) Tuning Est. S.E. Cov. Est. S.E. Cov. Est. S.E. Cov.
-0.25 -0.10 0.25 -0.10 0.09 99% -0.10 0.05 94% -0.12 0.07 96%
-0.25 -0.10 0.50 -0.08 0.06 99% -0.10 0.04 94% -0.11 0.05 98%
-0.25 -0.10 1.00 -0.06 0.04 97% -0.08 0.03 87% -0.09 0.03 96%
0.00 0.00 0.25 0.02 0.07 94% 0.00 0.03 94% -0.01 0.06 96%
0.00 0.00 0.50 0.03 0.06 98% 0.01 0.02 93% -0.00 0.05 97%
0.00 0.00 1.00 0.05 0.04 92% 0.03 0.02 80% 0.02 0.04 93%
0.25 0.32 0.25 0.36 0.15 98% 0.39 0.19 94% 0.36 0.17 94%
0.25 0.32 0.50 0.37 0.11 99% 0.38 0.14 95% 0.35 0.12 97%
0.25 0.32 1.00 0.34 0.08 99% 0.35 0.10 95% 0.33 0.10 98%
@@
Compared to results for lasso Riesz representers (LRR) [Chernozhukov et al., 2022c], KRRR considerably reduces the absolute bias in this smooth design. See [Chernozhukov et al., 2023, Tables S4-S6] for values analogous to those in Table 2. In particular at the location $v^{*}=0.25$, for lasso regression, KRRR gives absolute bias $(0.06,0.05,0.03)$ while LRR gives absolute bias $(0.09,0.13,0.11)$; for random forest regression, KRRR gives $(0.04,0.04,0.02)$ while LRR gives $(0.05,0.08,0.07)$; and for neural network regression, KRRR gives $(0.02,0.02,0.00)$ while LRR gives $(0.07,0.07,0.06)$.
7.2 Inferring heterogeneous treatment effects of 401(k) eligibility by age
Having validated the pointwise coverage of KRRR for heterogeneous treatment effects, I now use the procedure to infer the heterogeneous treatment effects of 401(k) eligibility on assets by age. The empirical analysis yields meaningful economic insights.
I follow the identification strategy of [Poterba et al., 1995]. The authors assume that when 401(k) was introduced, workers ignored whether jobs offered 401(k) plans and instead made employment decisions based on income and other observable job characteristics. They assume 401(k) eligibility was as good as randomly assigned, conditional on these covariates.
@@ 7.2 Inferring heterogeneous treatment effects of 401(k) eligibility by age
As in the simulations, I consider low and high dimensional specifications, and various choices of the nonparametric regression estimator $\hat{\gamma}$: lasso, random forest, and neural network. I use KRRR $\tilde{\alpha}$ with the Gaussian kernel. See Appendix E for implementation details.
Figures 2 and 3 visualize point estimates and pointwise 95% confidence intervals based on KRRR, in black. I find that 401(k) eligibility has positive and statistically significant effects that vary by age. The effects appear to be small and positive for 30 year olds (about \$2,500), and generally increasing, before plateauing for 45 to 60 year olds (about \$12,500). The effect is not statistically significant for 65 year olds.
In summary, under the stated assumptions, 401(k) eligibility seems to cause middle aged individuals to save about five times more than 30 years olds. Individuals in the former subpopulation tend to be at the peak of their earning potentials, while those in the latter subpopulation are closer to the beginning of their careers. The results are robust, across variations of the specification and the regression estimator.
For comparison, I also visualize the smooth nonparametric estimator of [Singh et al., 2024], in grey. The confidence intervals of this work corroborate their estimate, and quantify uncertainty. In particular, this paper’s confidence intervals suggest that the apparent dip in effects for 65 year olds is not statistically significant.
Figure 2: Heterogeneous treatment effects by age: Low dimensional specification. The point estimates and confidence intervals are constructed using KRRR. The smooth curve is a comparison estimator [Singh et al., 2024].
Figure 3: Heterogeneous treatment effects by age: High dimensional specification. The point estimates and confidence intervals are constructed using KRRR. The smooth curve is a comparison estimator [Singh et al., 2024].
@@ 8 Discussion
Causal inference introduces a counterfactual effective dimension. Under standard regularity conditions on causal parameters from semiparametric theory, e.g. a propensity score bounded away from zero and one, I show that the counterfactual effective dimension is not much greater than the actual effective dimension. Using this technique, I prove population $L_{2}$ rates of generalization error for kernel ridge balancing weights, similar to kernel ridge regression. These rates connect kernel ridge balancing weights with general semiparametric results which tolerate $\gamma_{0}\not\in H$ and which provide inference for causal functions.
@@
Appendix A Inference on causal functions
@@ A.1 Framework
I provide further details for the pointwise approximation of nonparametric causal functions. I discuss parameters of the form
@@
Recall Example 3.2. Here, $W=(D,X)$ concatenates the treatment $D\in\{0,1\}$ and covariates $X$, and we wish to study heterogeneity with respect to the first covariate $X_{1}\in\mathbb{R}$. Let $x_{1}^{*}$ be a particular value of $X_{1}$. Then the heterogeneous treatment effect for the subpopulation with $X_{1}=x_{1}^{*}$ is $\theta_{0}^{\lim}=\lim_{h\rightarrow 0}\theta_{0}^{h}$ where $\theta_{0}^{h}=\mathbb{E}\{\ell_{h}(W)\{\gamma_{0}(1,X)-\gamma_{0}(0,X)\}$,
@@ A.2 Technical details
It is convenient to refer to the Riesz representer of $f\mapsto\mathbb{E}\{m_{h}(W,f)\}$ by $\alpha_{h}$ and the Riesz representer of $f\mapsto\mathbb{E}\{\check{m}(W,f)\}$ by $\check{\alpha}_{0}$. In Example 3.2,
@@
Lemma 5.2 in the main text holds for $m_{h}$ and $\alpha_{h}$. In particular the rate conditions are for $\mathcal{R}(\hat{\alpha}_{h})$, i.e. the rate at which $\hat{\alpha}_{h}$ converges to $\alpha_{h}$.
I now clarify how various quantities in Lemma 5.2 depend on $h$. Consider the estimator $\hat{\alpha}_{h}(W)=\ell_{h}(W)\tilde{\alpha}(W)$, where $\tilde{\alpha}$ is the KRRR estimator for $\check{\alpha}_{0}$. Section 7 confirms that this choice performs well in simulations. The results below are in terms of $h$ and $\mathcal{R}(\tilde{\alpha})=\mathbb{E}[\{\tilde{\alpha}(W)-\check{\alpha}_{0}(W)\}^{2}|(W_{i})_{i\in[n]}]$, i.e. the rate at which $\tilde{\alpha}$ converges to $\check{\alpha}_{0}$.
Lemma A.1 (Characterization of key quantities; Theorem 2 of [Chernozhukov et al., 2023]).
@@
The other key quantities in Lemma 5.2 satisfy
@@
Corollary A.1 (Inference for causal functions; Corollary S1 of [Chernozhukov et al., 2023]).
Suppose the conditions of Lemmas 5.2 and A.1 hold as well as correct specification. As $n\rightarrow\infty$, suppose the bandwidth satisfies $h=o(1)$ and $n^{-1/2}h^{-3/2}=o(1)$. Suppose the following learning rate conditions hold:

### v4

@@ 6   Proof via the counterfactual effective dimension
The steps are as follows: (i) translate the desired statement from generalization error to $H$ norm (Lemma 8); (ii) characterize high probability events via concentration in $H$ (Proposition 4); (iii) use these high probability events to prove a nonasymptotic bound in terms of standard learning theory quantities (Proposition 5).
Among the high probability events is one that involves the counterfactual features $\phi^{(m)}(W)$. It pertains to the variance of the KRRR estimator $\tilde{\alpha}$. To justify this high probability event, I bound the counterfactual effective dimension of $\phi^{(m)}(W)$ in terms of the actual effective dimension of $\phi(W)$ (Lemma 9).
The main result (Theorem 2) then follows from simplifying the nonasymptotic bound (Proposition 5) using known bounds on the learning theory quantities that hold under smoothness (Assumption 2) and spectral decay (Assumption 3) conditions. The latter controls the actual effective dimension. I defer proofs of lemmas to Appendix D.
Lemma 8 (From generalization error to $H$ norm; c.f. Lemma 12 of [Fischer and Steinwart, 2020]).
Suppose the conditions of Lemma 3 hold. Then $\mathcal{R}(\tilde{\alpha})=\|T^{\frac{1}{2}}(\tilde{\alpha}-\alpha_{H})\|^{2}_{H}.$
Definition 5 (Standard learning theory quantities).
@@
Definition 6 (New learning theory quantity).
Define the counterfactual effective dimension $\mathcal{N}^{(m)}(\lambda)=Tr\{(T+\lambda)^{-1}T^{(m)}\}$, where $T^{(m)}=\mathbb{E}[\phi^{(m)}(W)\otimes\{\phi^{(m)}(W)\}^{*}]$ is the counterfactual covariance operator.
Lemma 9 (New learning theory bound).
Suppose the conditions of Lemma 3 hold. Then $\mathcal{N}^{(m)}(\lambda)\leq\bar{M}\cdot\mathcal{N}(\lambda).$
Remark 5 (Avoiding an auxiliary approximation assumption).
Lemma 9 demonstrates, by a simple argument, that a standard assumption in semiparametric theory (Assumption 1) implies that the counterfactual effective dimension (Definition 6) is not much greater than the actual effective dimension (Definition 5). This connection between semiparametric theory and learning theory appears to be new, and it powers the main result.
Previous works that study generalization error, for different estimators of Riesz representers, appear to place auxiliary approximation assumptions to limit the counterfactual effective dimension, e.g. [Chernozhukov et al., 2022b, Assumption SC(a)], [Chernozhukov et al., 2022a, Assumption 2], and [Chernozhukov et al., 2020, Assumption 2].
Definition 7 (High probability events in $H$).
@@
Proposition 4 (High probability events in $H$).
Suppose the conditions of Lemma 3 hold. Then $\mathbb{P}(\mathcal{E}_{j}^{c})\leq\delta/3$ for each $j\in\{1,2,3\}.$
The proof of Proposition 4 appeals to Lemma 9 in order to handle $\mathcal{E}_{3}$, which contains the counterfactual features via the counterfactual mean embedding $\hat{\mu}^{(m)}=\mathbb{E}_{n}\{\phi^{(m)}(W)\}$.
Proposition 5 (Nonasymptotic bound).
Suppose the conditions of Lemma 3 hold. If $n$ is sufficiently large that $n\geq 192\ln^{2}(6/\delta)\kappa^{\prime}\mathcal{N}(\lambda)\lambda^{-1}$ and $\lambda\leq\|T\|_{op},$ then with probability $1-\delta$, $\|T^{1/2}(\tilde{\alpha}-\alpha_{H})\|^{2}_{H}\leq 96\ln^{2}(6/\delta)\mathopen{}\mathclose{{}\left\{\mathcal{A}(\lambda)+\frac{\kappa^{2}\mathcal{B}(\lambda)}{n^{2}\lambda}+\frac{\kappa\mathcal{A}(\lambda)}{n\lambda}+\frac{\kappa^{\prime}\Upsilon^{2}}{n^{2}\lambda}+\frac{\Sigma^{2}\mathcal{N}(\lambda)}{n}}\right\}.$
The proof of Proposition 5 appeals to Proposition 4 and the union bound.
@@ 7   Simulated and real data analysis
7.1   Nominal pointwise coverage of heterogeneous effects
I present coverage simulations for heterogeneous treatment effects (Example 2) viewed as a causal function. To lighten notation, I let $V=X_{1}$ and I write the heterogeneous treatment effects as $\textsc{cate}(v)$.
@@ 7.1   Nominal pointwise coverage of heterogeneous effects
Figure 1: Simulation design. cate(v) is the curve and values of $\theta_{0}$ are the three points.
I evaluate confidence sets constructed from the KRRR estimator $\tilde{\alpha}$ (Definition 2) in the context of debiased machine learning $(\hat{\theta},\hat{\sigma}^{2})$ (Definition 4). I consider three variations of $(\hat{\theta},\hat{\sigma}^{2})$, corresponding to different choices of the nonparametric regression estimator $\hat{\gamma}$: lasso, random forest, and neural network. Previous works on classical kernel ridge balancing weights typically disallow $\gamma_{0}\not\in H$, and appear not to justify inference on heterogeneous treatment effects. This use of kernel ridge balancing weights is justified by the main result of this paper.
Following [Chernozhukov et al., 2023, Tables S1 to S6], I consider low dimensional and high dimensional specifications. In the low dimensional specification, the estimator $\hat{\gamma}$ uses $(D,X)$ as well as their interactions. In the high dimensional specification, $\hat{\gamma}$ uses fourth order polynomials. I focus here on the latter. See Appendix E for the former.
Throughout, KRRR uses the Gaussian kernel for $X$, which corresponds to using all Hermite polynomials, appropriately weighted. See Appendix E for implementation details.
An important choice is how to tune the bandwidth $h$ in $\textsc{local}\{(v-v^{*})/h\}$. I follow the heuristic $h=c_{h}\hat{\sigma}_{v}n^{-1/5}$ of previous work, where $\hat{\sigma}^{2}_{v}$ is the sample variance of $V$. This heuristic satisfies the theoretical requirements $h=o(1)$ and $n^{-1/2}h^{-3/2}=o(1)$, and it implies a nonparametric rate of convergence of $(nh)^{-1/2}=n^{-2/5}$. The hyperparameter $c_{h}$ is chosen by the analyst. To evaluate robustness to tuning, I consider $c_{h}\in\{0.25,0.50,1.00\}$.
@@
Compared to results for lasso Riesz representers (LRR) [Chernozhukov et al., 2022a], KRRR considerably reduces the absolute bias in this smooth design. See [Chernozhukov et al., 2023, Tables S4-S6] for values analogous to those in Table 1. In particular at the location $v^{*}=0.25$, for lasso regression, KRRR gives absolute bias $(0.06,0.05,0.03)$ while LRR gives absolute bias $(0.09,0.13,0.11)$; for random forest regression, KRRR gives $(0.04,0.04,0.02)$ while LRR gives $(0.05,0.08,0.07)$; and for neural network regression, KRRR gives $(0.02,0.02,0.00)$ while LRR gives $(0.07,0.07,0.06)$.
7.2   Confidence sets for heterogeneous effects of 401(k)
Having justified the pointwise coverage of KRRR for heterogeneous treatment effects, I now use the procedure to quantify uncertainty for the heterogeneous effects of 401(k) eligibility on assets by age. The empirical analysis yields meaningful economic insights.
I follow the identification strategy of [Poterba et al., 1995]. The authors assume that when 401(k) was introduced, workers ignored whether jobs offered 401(k) plans and made employment decisions based on income and other observable job characteristics. They assume 401(k) eligibility was as good as randomly assigned, conditional on covariates.
@@ 7.2   Confidence sets for heterogeneous effects of 401(k)
As in the simulations, I consider a high dimensional specification and various choices of the nonparametric regression estimator $\hat{\gamma}$: lasso, random forest, and neural network. I use KRRR $\tilde{\alpha}$ with the Gaussian kernel. See Appendix E for the low dimensional specification.
Figures 2 and 3 visualize point estimates and pointwise 95% confidence sets based on KRRR, in black. I find that 401(k) eligibility has positive and statistically significant effects that vary by age. The effects appear to be small and positive for 30 year olds (about \$2,500), and generally increasing, before plateauing for 45 to 60 year olds (about \$12,500). The effect is not statistically significant for 65 year olds.
In summary, under the stated assumptions, 401(k) eligibility seems to cause middle aged individuals to save about five times more than 30 years olds. Individuals in the former subpopulation tend to be at the peak of their earning potentials, while those in the latter subpopulation are closer to the beginning of their careers. The results are robust, across variations of the regression estimator and the specification.
For comparison, I also visualize the smooth nonparametric estimator of [Singh et al., 2024], in grey. The confidence sets of this work corroborate their estimate, and quantify uncertainty. In particular, this paper’s confidence sets suggest that the apparent dip in effects for 65 year olds is not statistically significant.
Figure 2: Heterogeneous treatment effects by age. The point estimates and confidence sets use KRRR. The smooth curve is a comparison estimator [Singh et al., 2024].
@@ 8   Discussion
Causal inference introduces a counterfactual effective dimension. Under standard regularity conditions on causal parameters from semiparametric theory, e.g. a propensity score bounded away from zero and one, I show that the counterfactual effective dimension is not much greater than the actual effective dimension. Using this technique, I prove population $L_{2}$ rates of generalization error for kernel ridge balancing weights, similar to kernel ridge regression. These rates connect kernel ridge balancing weights with general semiparametric results which tolerate $\gamma_{0}\not\in H$ and justify confidence sets for causal functions.
@@
Appendix A Formal inference on causal functions
@@ A.1   Framework
By formal inference, I mean justification for Gaussian approximation and nominal coverage guarantees. In this section, I provide further details for the pointwise approximation of nonparametric causal functions. I discuss parameters of the form
@@
Recall Example 2. Here, $W=(D,X)$ concatenates the treatment $D\in\{0,1\}$ and covariates $X$, and we wish to study heterogeneity with respect to the first covariate $X_{1}\in\mathbb{R}$. Let $x_{1}^{*}$ be a particular value of $X_{1}$. Then the heterogeneous treatment effect for the subpopulation with $X_{1}=x_{1}^{*}$ is $\theta_{0}^{\lim}=\lim_{h\rightarrow 0}\theta_{0}^{h}$ where $\theta_{0}^{h}=\mathbb{E}\{\ell_{h}(W)\{\gamma_{0}(1,X)-\gamma_{0}(0,X)\}$,
@@ A.2   Technical details
It is convenient to refer to the Riesz representer of $f\mapsto\mathbb{E}\{m_{h}(W,f)\}$ by $\alpha_{h}$ and the Riesz representer of $f\mapsto\mathbb{E}\{\check{m}(W,f)\}$ by $\check{\alpha}_{0}$. In Example 2,
@@
Lemma 7 in the main text holds for $m_{h}$ and $\alpha_{h}$. In particular the rate conditions are for $\mathcal{R}(\hat{\alpha}_{h})$, i.e. the rate at which $\hat{\alpha}_{h}$ converges to $\alpha_{h}$.
I now clarify how various quantities in Lemma 7 depend on $h$. Consider the estimator $\hat{\alpha}_{h}(W)=\ell_{h}(W)\tilde{\alpha}(W)$, where $\tilde{\alpha}$ is the KRRR estimator for $\check{\alpha}_{0}$. Section 7 confirms that this choice performs well in simulations. The results below are in terms of $h$ and $\mathcal{R}(\tilde{\alpha})=\mathbb{E}[\{\tilde{\alpha}(W)-\check{\alpha}_{0}(W)\}^{2}|(W_{i})_{i\in[n]}]$, i.e. the rate at which $\tilde{\alpha}$ converges to $\check{\alpha}_{0}$.
Lemma 10 (Characterization of key quantities; Theorem 2 of [Chernozhukov et al., 2023]).
@@
The other key quantities in Lemma 7 satisfy
@@
Corollary 9 (Inference for causal functions; Corollary S1 of [Chernozhukov et al., 2023]).
Suppose the conditions of Lemmas 7 and 10 hold as well as correct specification. As $n\rightarrow\infty$, suppose the bandwidth satisfies $h=o(1)$ and $n^{-1/2}h^{-3/2}=o(1)$. Suppose the following learning rate conditions hold: